# Cross-Lingual F5-TTS: Towards Language-Agnostic Voice Cloning and Speech Synthesis

### v2

@@ 1 Introduction
Among NAR models, flow-matching-based approaches have emerged as particularly promising. For example, VoiceBox[7] surpasses AR-based VALL-E[1] in Word Error Rate (WER) and Speech Similarity (SIM) with 20 times faster inference speed. E2 TTS[8] and F5-TTS[9] have achieved remarkable improvements with their simpler architectures. Recent cross-lingual TTS approaches, such as VALL-E X [13], have achieved impressive cross-lingual voice cloning capabilities by training on large-scale multilingual data. However, all these models still rely on audio prompt transcripts, which creates inherent challenges when reference transcripts are unavailable. Different flow-matching-based models employ various duration modeling methods. While some flow-matching-based models use explicit phoneme duration predictors[7] , F5-TTS estimates duration by the ratio of the audio prompt transcript length and target text length. However, this mechanism breaks down in a cross-lingual context, as text-length ratios between different languages may not correspond to speech duration ratios.
In this paper, we propose Cross-Lingual F5-TTS, a novel framework based on F5-TTS that achieves cross-lingual zero-shot voice cloning by dropping audio prompt transcripts. Leveraging the MMS[14] forced alignment, we preprocess our training data to identify word boundaries. To solve the challenge of duration prediction, we introduce dedicated speaking rate predictors at three linguistic granularities including phoneme, syllable, and word levels. These models are trained to estimate speaking rate from the audio prompt, providing a robust, language-independent mechanism for determining the duration of the target utterance. Our approach achieves comparable performance to F5-TTS on LibriSpeech-PC test-clean and Seed-TTS-eval, while successfully extending its capabilities to cross-lingual scenarios with promising results.  [^1] An online demo can be found at https://qingyuliu0521.github.io/Cross_lingual-F5-TTS/. The model will be open-sourced on Hugging Face.
@@ 3.4 Evaluation
4 Results
4.1 Speaking Rate Predictor Performance

### v3

@@ 1 Introduction
Among NAR models, flow-matching-based approaches have emerged as particularly promising. For example, VoiceBox[7] surpasses AR-based VALL-E[1] in Word Error Rate (WER) and Speech Similarity (SIM) with 20 times faster inference speed. E2 TTS[8] and F5-TTS[9] have achieved remarkable improvements with their simpler architectures. Recent cross-lingual TTS approaches, such as VALL-E X [13], have achieved impressive cross-lingual voice cloning capabilities by training on large-scale multilingual data. However, all these models still rely on audio prompt transcripts, which creates inherent challenges when reference transcripts are unavailable. Different flow-matching-based models also employ various duration modeling methods. While some flow-matching-based models use explicit phoneme duration predictors[7] , F5-TTS estimates duration by the ratio of the audio prompt transcript length and target text length. However, this mechanism breaks down in a cross-lingual context, as text-length ratios between different languages may not correspond to speech duration ratios.
In this paper, we propose Cross-Lingual F5-TTS, a novel framework based on F5-TTS that achieves cross-lingual zero-shot voice cloning by dropping audio prompt transcripts. Leveraging the MMS[14] forced alignment, we preprocess our training data to identify word boundaries. To solve the challenge of duration prediction, we introduce dedicated speaking rate predictors at three linguistic granularities including phoneme, syllable, and word levels. These models are trained to estimate speaking rate from the audio prompt, providing a robust, language-independent mechanism for determining the duration of the target utterance. Our approach achieves comparable performance to F5-TTS on LibriSpeech-PC test-clean and Seed-TTS-eval, while successfully extending its capabilities to cross-lingual scenarios with promising results.  [^1] Audio samples can be found at https://qingyuliu0521.github.io/Cross_lingual-F5-TTS/. The interactive demo is available at https://huggingface.co/spaces/chenxie95/Cross-Lingual_F5-TTS_Space.
@@ 4.1 Speaking Rate Predictor Performance
4.2 Intra-lingual Voice Cloning Performance
@@ 5 Conclusion
6 Acknowledgments
This work was supported by the Science and Technology Innovation (STI) 2030-Major Project (2022ZD0208700), National Natural Science Foundation of China (No. U23B2018), Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0102 and Yangtze River Delta Science and Technology Innovation Community Joint Research Project (2024CSJGG011 00).